Find the following limit.:


Lim x →0 for 1-cos x²/x².sin x²

1
Expert's answer
2022-03-20T06:44:34-0400

"\lim\limits _{x\to 0}\frac{1-cosx^2}{x^2sinx^2}=\\
\colorbox{aqua}{We use formulas}\\
1-cosa=2sin^2\frac{a}{2}\\
sina=2sin\frac{a}{2}cos\frac{a}{2}\\
=\lim\limits _{x\to 0}\frac{2sin^2\frac{x^2}{2}}{x^22sin\frac{x^2}{2}cos\frac{x^2}{2}}=\lim\limits _{x\to 0}\frac{sin\frac{x^2}{2}}{x^2cos\frac{x^2}{2}}=\\
\colorbox{aqua}{We use formulas}\\
\lim\limits _{x\to 0}\frac{sinx}{x}=1\\

=\lim\limits _{x\to 0}\frac{sin\frac{x^2}{2}}{2\cdot\frac{x^2}{2}cos\frac{x^2}{2}}=
\frac{1}{2cos0}=\frac{1}{2}"